OpenStudy (anonymous):

It was reported that about 80% of airline tickets-nearly $65 billion worth last year are issued over the internet. What were the total sales of airline tickets last year? Round to the nearest billion.

OpenStudy (anonymous):

same thing as the last use a proportion 65/80%=x/100%. x=81.25

OpenStudy (anonymous):

thanks again.. so much